how do you solve the polynomial equation 3x^2+11x-4=0
Answer by shirsenduban(10)

3x^2+11x-4=0
3x^2+12x-x-4=0
3x(x+4)-1(x+4)=0
(3x-1)(x+4)=0
3x=1 or x=-4
x=1/3 or -4